# Simulation-based inference with scattering representations: scattering is all you need

### v1

@@ Abstract
We demonstrate the first successful use of scattering representations without further compression for simulation-based inference (SBI) with images (i.e. field-level), illustrated with a cosmological case study. Scattering representations provide a highly effective representational space for subsequent learning tasks, although the higher dimensional compressed space introduces challenges. We overcome these through spatial averaging, coupled with more expressive density estimators. Compared to alternative methods, such an approach does not require additional simulations for either training or computing derivatives, is interpretable, and resilient to covariate shift. As expected, we show that a scattering only approach extracts more information than traditional second order summary statistics.
@@ 1 Introduction
For image data, one such approach is through the use of higher order statistics [e.g. 47, 13, 25, 38, 39, 40, 41]. These statistics, however, are difficult to compute and it is often even more difficult to accurately characterise their covariance structure for use with an analytical likelihood. Another direction involves the use of image-level (commonly called field-level) inference [e.g. 45, 9, 49, 6, 46]. Such approaches, however, must overcome the challenges of high dimensional Markov chain Monte Carlo (MCMC) sampling and approximate likelihoods. SBI has recently been demonstrated as an effective alternative for field-level inference, eliminating these challenges [27, 22, 31]. However, due to the high dimensionality of field-level data, there is a need for substantial data compression.
@@
Wavelet scattering representations provide an alternative summary statistic [34, 12]. Whilst scattering representations are inspired by CNNs, they are designed and not learnt. This has the distinct advantage of not being a neural method. Consequently, they are interpretable [17, 2, 16] and do not require any training data. Scattering representations have gained traction in cosmology recently and appear to be highly effective at capturing informative non-Gaussian information [17, 2, 16, 50, 18]. In the asymptotic case of infinite depth they capture all information content of the field analysed, while also capturing the majority of information in a relatively small number of levels [34, 36, 16]. Furthermore, scattering transforms have recently been considered in cosmology as a form of compression for SBI [22, 8]. However, they have either been combined with neural compression [22] or demonstrated to provide unstable inferences [8]. The use of scattering representations without further neural compression has not been demonstrated successfully.
@@ 5 Conclusions
We demonstrate for the first time that field-level SBI can be performed effectively using wavelet scattering representations alone to provide a compressed representation without any further neural compression. The challenge of a higher dimensional compressed representation is overcome by incorporating spatial averaging and a judicious choice of parameters, coupled with more expressive density estimators. This approach has considerable advantages compared to the alternative statistical or neural compression approaches typically considered in that it does not require additional simulations to train a neural compressor, it does not require simulation suites with perturbed parameters needed to compute numerical derivatives for statistical compression approaches, it is more interpretable than the use of black-box neural networks, and it is not sensitive to covariate shift. Furthermore, we demonstrate that SBI with scattering representations extracts beyond two-point statistical information, as expected, since recovered parameter constraints are distinct and tighter than those recovered by bandpowers (thus, scattering representations and bandpowers can be combined to improve constraining power further still). In a sense, for field-level SBI, scattering is all you need for compression. That said, we do not dogmatically advocate that scattering alone should always be used. While the scattering approach presented has numerous advantages, our findings motivate further studies directly comparing the effectiveness of different compression approaches for field-level SBI.

### v2

@@ Abstract
We demonstrate the successful use of scattering representations without further compression for simulation-based inference (SBI) with images (i.e. field-level), illustrated with a cosmological case study. Scattering representations provide a highly effective representational space for subsequent learning tasks, although the higher dimensional compressed space introduces challenges. We overcome these through spatial averaging, coupled with more expressive density estimators. Compared to alternative methods, such an approach does not require additional simulations for either training or computing derivatives, is interpretable, and resilient to covariate shift. As expected, we show that a scattering only approach extracts more information than traditional second order summary statistics.
@@ 1 Introduction
For image data, one such approach is through the use of higher order statistics [e.g. 48, 13, 25, 38, 39, 40, 41, 29]. These statistics, however, are difficult to compute and it is often even more difficult to accurately characterise their covariance structure for use with an analytical likelihood. Another direction involves the use of image-level (commonly called field-level) inference [e.g. 46, 9, 50, 6, 47, 42, 29]. Such approaches, however, must overcome the challenges of high dimensional Markov chain Monte Carlo (MCMC) sampling and approximate likelihoods. SBI has recently been demonstrated as an effective alternative for field-level inference, eliminating these challenges [27, 22, 32]. However, due to the high dimensionality of field-level data, there is a need for substantial data compression.
@@
Wavelet scattering representations provide an alternative summary statistic [35, 12]. Whilst scattering representations are inspired by CNNs, they are designed and not learnt. This has the distinct advantage of not being a neural method. Consequently, they are interpretable [17, 2, 16] and do not require any training data. Scattering representations have gained traction in cosmology recently and appear to be highly effective at capturing informative non-Gaussian information [17, 2, 16, 51, 18]. In the asymptotic case of infinite depth they capture all information content of the field analysed, while also capturing the majority of information in a relatively small number of levels [35, 37, 16]. Furthermore, scattering transforms have recently been considered in cosmology as a form of compression for SBI [22, 8]. However, they have either been combined with neural compression [22] or demonstrated to provide unstable inferences [8].
@@ 5 Conclusions
We demonstrate that field-level SBI can be performed effectively using wavelet scattering representations alone to provide a compressed representation without any further neural compression. The challenge of a higher dimensional compressed representation is overcome by incorporating spatial averaging and a judicious choice of parameters, coupled with more expressive density estimators. This approach has considerable advantages compared to the alternative statistical or neural compression approaches typically considered in that it does not require additional simulations to train a neural compressor, it does not require simulation suites with perturbed parameters needed to compute numerical derivatives for statistical compression approaches, it is more interpretable than the use of black-box neural networks, and it is not sensitive to covariate shift. Furthermore, we demonstrate that SBI with scattering representations extracts beyond two-point statistical information, as expected, since recovered parameter constraints are distinct and tighter than those recovered by bandpowers (thus, scattering representations and bandpowers can be combined to improve constraining power further still). In parallel with this work, successful use of scattering-only compression for SBI has simultaneously been demonstrated for 21cm epoch of reionization signals [54], further evidencing the effectiveness of scattering for SBI. In a sense, for field-level SBI, scattering is all you need for compression. That said, we do not dogmatically advocate that scattering alone should always be used. While the scattering approach presented has numerous advantages, our findings motivate further studies directly comparing the effectiveness of different compression approaches for field-level SBI.